OpenStudy (anonymous):

2sec^2(x)+tan^2(x)-3=0.....lost

OpenStudy (cgreenwade2000):

There is a substitution for sec^2 of x that we should know here.

OpenStudy (ajprincess):

\(sec^2x=1+tan^2x\)

OpenStudy (ajprincess):

Can u solve it nw @jotwestmoreland?

OpenStudy (ajprincess):

Substitute (\1+tan^2x\) in place of \(sec^2x\) and try solving. can u do @jotwestmoreland?

OpenStudy (anonymous):

@ajprincess i get stuck at 2+2tan^2x+tan^2x-3=0

OpenStudy (ajprincess):

\[2+2tan^2x+tan^2x-3=0\] \[3tan^2x-1=0\] \[3tan^2x=1\] \[tan^2x=\frac{1}{3}\] \[tanx=\frac{1}{\sqrt3}\] \[x=tan^{-1}(\frac{1}{\sqrt3})\] \[x=30^o\] Is it clear @jotwestmoreland?